OpenStudy (anonymous):

what is 1 1/8 divided by 1/3

OpenStudy (anonymous):

33/8

OpenStudy (anonymous):

Is that 11 divided by 8 or 1 and 1/8?

OpenStudy (anonymous):

11/8*3 = 33/8

OpenStudy (anonymous):

1 & 1/8 sorry

OpenStudy (anonymous):

divide by 1/3 means multiply by 3

OpenStudy (anonymous):

First convert the mixed numeral to an improper fraction. so 1 1/8 = 9/8. Then, multiply this by the reciprocal of the "divided by" part. (9/8)*(3/1) = 27/8

OpenStudy (nikita2):

3 and 3/8 )